to the right of z= -1.89


1
Expert's answer
2021-05-18T15:13:26-0400
"P(Z>-1.89)=1-P(Z\leq-1.89)"

"\approx1-0.0293790\approx0.970621"

By the symmetry of the bell curve, finding the area to the right of a negative 

z-score is equivalent to the area to the left of the corresponding positive 

z-score.


"P(Z>-1.89)=P(Z<1.89)"

"\approx0.970621"



"P(Z>-1.89)=0.970621"